OpenStudy (anonymous):

Find the limit of the function algebraically. limit as x -->9 of quantity x^2-81/x-9 .

OpenStudy (anonymous):

\(x^2-81\) is a difference of squares... consider \(x^2-81=x^2-9x+9x-81=x(x-9)+9(x-9)=(x-9)(x+9)\)

OpenStudy (anonymous):

THANKS @oldrin.bataku